OpenStudy (beccaboo333):

Write the equation of the line that passes through (-3, 4) with a slope of 2 in point-slope form.

OpenStudy (beccaboo333):

A. y - 4 = 2(x + 3) B. y + 4 = 2(x - 3) C. y + 3 = 2(x - 4) D. y - 3 = 2(x + 4)

OpenStudy (anonymous):

the equation of a line is: y=mx +b M = slope your ordered pairs are (x,y) plug your values in the formula and there you have it.

OpenStudy (anonymous):

oh point slope form

OpenStudy (beccaboo333):

I don't understand it like i had said to ganeshie8 before.. this stuff is still pretty confusing

OpenStudy (beccaboo333):

help please?

ganeshie8 (ganeshie8):

point slope form :- \(\large y - y_1 = m(x-x_1) \)

ganeshie8 (ganeshie8):

you need to memorize above equation for line for the rest of time of ur course in Algebra :)

ganeshie8 (ganeshie8):

we're given :- point (-3, 4) \(x_1\) \(y_1\)

ganeshie8 (ganeshie8):

ALso, we're given slope = \(m\) = 2

ganeshie8 (ganeshie8):

plug them in above equation for line, wat do u get ?

ganeshie8 (ganeshie8):

\(\large y - y_1 = m(x-x_1)\) \(\large y - 4 = 2(x- - 3)\) \(\large y - 4 = 2(x + 3)\)

ganeshie8 (ganeshie8):

thats ur answer, i want u spend sometime digest a bit :)

OpenStudy (beccaboo333):

o.O this stuff is gonna take me a while to remember. thanks again. 5 other questions.. still willing to help if i cant figure them out?

ganeshie8 (ganeshie8):

il try, close this and post a new q :)